OpenStudy (anonymous):

Help please, how do I find the inverse of the following: f(x)=(In(2-x))/(x^2+3x), h(t)= Sin^-1(t-4) and g(x)=(e^(sinx))/In(x-3)

OpenStudy (anonymous):

I meant how do I find the domain of the following functions?

OpenStudy (helder_edwin):

for the first one u have to solve: \[ \large 2-x>0 \] and \[ \large x^2+3x=0 \] excluding whatever u obtain from this last.

OpenStudy (anonymous):

domain of arcsine is \([-1,1]\) so for the second one solve \(-1\leq t-4\leq 1\) for \(t\)

OpenStudy (anonymous):

last one is trickier domain of log is \((0,\infty)\) so you have to solve \(x-3>0\) but you also have to make sure that \(\ln(x-3)\neq 0\) because you cannot divide by zero

OpenStudy (anonymous):

\(\ln(1)=0\) so make sure \(x-3\neq 1\)

OpenStudy (anonymous):

Hey Satellite, should i just use the domain of sinx as the domain for the denominator?

OpenStudy (anonymous):

which one?

OpenStudy (anonymous):

for the last one

OpenStudy (anonymous):

\[ g(x)=\frac{e^{sinx}}{\ln(x-3)}\]

OpenStudy (anonymous):

sine has no restriction, neither does \(e^x\) so no worries in the numerator you only have to worry about the denominator

OpenStudy (anonymous):

you have to make sure that the log is defined, so \(x-3>0\iff x>3\)

OpenStudy (anonymous):

also you have to make sure that \(\ln(x-3)\neq 0\) because you cannot divide by 0

OpenStudy (anonymous):

therefore \(x-3\neq 1\iff x\neq 4\)

OpenStudy (anonymous):

so your "final answer" would be \(x>3,x\neq 4\) however you would like to write it,

OpenStudy (anonymous):

Thanks a lot Satellite, you really helped. Smile today!!!!

OpenStudy (anonymous):

ok i will